OpenStudy (anonymous):

solve a linear equation: 6=3/2x+1/2(x-4)

OpenStudy (anonymous):

x=8-sqrt61 and 8+sqrt61

OpenStudy (anonymous):

wait what???

Directrix (directrix):

solve a linear equation ------------------ 6= (3/2) x+(1/2) (x-4) 6 = 2x -2 x = 4

Directrix (directrix):

Linear equations don't usually have irrational solutions.

OpenStudy (anonymous):

how did you get rid of the fractions

Directrix (directrix):

6= (3/2) x+(1/2) (x-4) Multiply through by 2 --> Multipllication property of Equality 2*6 = 2 * (3/2) x + 2 * (1/2) ( x - 4) 12 = 3x + ( x - 4) 12 + 4 = 4x 16 = 4x 4 = x

Directrix (directrix):

Question?